Circular No. 6841

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998V IN NGC 6627 G. M. Hurst, Basingstoke, England, reports the discovery of an ap-

parent supernova (mag ∼ 16) by Mark Armstrong, Rolvenden, on an unfiltered CCD image obtained on Mar. 10.219 UT with a 0.20-m reflector in the course of the U.K. Nova/Supernova Patrol. An image obtained on Mar. 12.115 by Armstrong yields mag 15.5 and the following position: α = 18^{h}22^{m}37^{s}.40, δ = +15^{o}42′08′′.4 (equinox 2000.0), which is ∼ 30′′ northwest of the center of NGC 6627. The object is not present in the Vickers CCD atlas or the Palomar Sky Survey, nor does it appear on earlier U.K. patrol images obtained during 1997 July 28–Oct. 21. T. Boles, Wellingborough, England, imaged the galaxy on Mar. 12.156, also finding SN 1998V near mag 15.5 (unfiltered CCD). Hurst further reports that there is no star present in the USNO A1.0 catalogue to mag 19.

S. Nakano, Sumoto, Japan, forwards the following precise position end figures and magnitude for SN 1998V, measured by Y. Kushida from a CCD image taken on Mar. 13.75 UT by R. Kushida with a 0.40-m reflector at Yatsugatake South Base Observatory: 37^{s}.40, 07′′.7, V = 15.1.

SUPERNOVA 1998S IN NGC 3877 Y. L. Qiu, Beijing Astronomical Observatory (BAO), writes: “The

low-dispersion spectra taken with 2-m telescope at Xinglong Station by L. Cao, Y. L. Qiu, and Q. Y. Qiao on Mar. 7.7 and 12.7 UT show that the spectral evolution of SN 1998S is peculiar. The spectrum obtained on Mar. 12 displayed a blue continuum, superimposed by weak Balmer lines, Hα, Hβ, Hδ, and Hγ. The features near 467 nm (He II and permitted N III/C III, as reported on IAUC 6832) in the spectrum of Mar. 7 nearly disappeared by Mar. 12. The profiles of Hα and Hβ in the spectrum of Mar. 12 change to be asymmetric, and the blue wings become steeper. The equivalent widths of Hα and Hβ are −0.83 nm, and −0.24 nm, respectively, compared to −3.2 nm and −1.3 nm on Mar. 7. The full width at the base of Hα is 7.4 nm, implying an expansion velocity of 3380 km/s.”

Visual magnitude estimates: Mar. 12.75 UT, 12.5 (L. Kiss, Szeged, Hungary); 13.22, 12.2 (R. Keen, Mt. Thorodin, CO; T UMa AAVSO chart); 14.799, 12.1 (K. Hornoch, Lelekovice, Czech Republic; GSC stars).

1998 March 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green